### v1

# HiSTF Mamba: Hierarchical Spatiotemporal Fusion with Multi-Granular Body-Spatial Modeling for High-Fidelity Text-to-Motion Generation

Abstract
Text-to-motion generation is a rapidly growing field at the nexus of multimodal learning and computer graphics, promising flexible and cost-effective applications in gaming, animation, robotics, and virtual reality. Existing approaches often rely on simple spatiotemporal stacking, which introduces feature redundancy, while subtle joint-level details remain overlooked from a spatial perspective. To this end, we propose a novel HiSTF Mamba framework. The framework is composed of three key modules: Dual-Spatial Mamba, Bi-Temporal Mamba, and Dynamic Spatiotemporal Fusion Module (DSFM). Dual-Spatial Mamba incorporates “Part-based + Whole-based” parallel modeling to represent both whole-body coordination and fine-grained joint dynamics. Bi-Temporal Mamba adopts a bidirectional scanning strategy, effectively encoding short-term motion details and long-term dependencies. DSFM further performs redundancy removal and extraction of complementary information for temporal features, then fuses them with spatial features, yielding an expressive spatio-temporal representation. Experimental results on the HumanML3D dataset demonstrate that HiSTF Mamba achieves state-of-the-art performance across multiple metrics. In particular, it reduces the FID score from 0.283 to 0.189, a relative decrease of nearly 30%. These findings validate the effectiveness of HiSTF Mamba in achieving high fidelity and strong semantic alignment in text-to-motion generation.
1 Introduction
Text-to-motion aims to generate human motion sequences based on natural language descriptions, and has emerged as a key research direction at the intersection of multimodal learning and computer graphics. Using textual input to drive human motion generation, text-to-motion provides a flexible and cost-effective approach for applications such as games [33], animation [27], robotic interaction [31], and virtual reality (VR).
Early motion generation has predominantly relied on two mainstream paradigms: Generative Adversarial Networks (GANs) [15] and Variational Autoencoders (VAEs) [28]. GAN-based approaches [32, 54, 11, 26, 5] leverage adversarial training to produce diverse motion sequences but suffer from mode collapse, leading to observable distortions in limb movement amplitudes [3]. Meanwhile, VAE-based models [6, 29, 55, 41] utilize latent-space representations to maintain temporal continuity; however, they often sacrifice motion details due to insufficient feature disentanglement. Although hybrid architectures [14] attempt to jointly capture spatio-temporal features, parameter redundancy and inadequate cross-dimensional interactions still limit their performance. In recent years, diffusion-based motion generation methods [24] have emerged rapidly. They employ a progressive generation mechanism grounded in probabilistic denoising to strike a balance between diversity and realism, thereby circumventing many inherent drawbacks of conventional GANs and VAEs. For example, MDM [51], MLD [9], MotionDiffuse [59], and MoFusion [12] have shown notable advantages in visual fidelity, motion smoothness, and scalability. To further enhance spatiotemporal modeling, FineMoGen [58] and Spatio-Temporal Graph Diffusion [36] design separate temporal and spatial modules, then adopt serial or parallel strategies to capture richer dependencies. Motion Mamba [61] applies Mamba [16] to these modules, improving efficiency and reducing training time.
However, despite the promising progress in efficiency, Motion Mamba[61] still has significant limitations in modeling fine-grained joint details and effectively fusing spatio-temporal features. The reasons for these limitations are two-fold. (1) Lacking the spatial modeling and feature learning of local body. Although feeding the entire motion matrix into the network at the same time can capture the overall motion pattern, its fixed information capacity severely compresses early joint data as training progresses, thereby neglecting subtle relationships between individual joints or limbs. Furthermore, Selective State Space Models (Selective SSMs) typically discards these essential local details, pushing the model toward broad motion structures rather than subtle joint dynamics. (2) Inadequate spatio-temporal fusion. While Motion Mamba attempts to improve the fusion of spatio-temporal features by stacking spatio-temporal modules, this approach often leads to repeated extraction of similar or overlapping features, leaving subsequent layers with limited benefit from redundancy and inefficient learning. These redundancies significantly increase computational costs, forcing the model to spend resources on removing duplicate features rather than polishing critical spatio-temporal details, thus reducing the fidelity of motion generation.
To address these limitations in modeling fine-grained joint details and effectively fusing spatio-temporal features, we propose HiSTF Mamba. HiSTF Mamba consists of three key components: Dual-Spatial Mamba, Bi-Temporal Mamba, and DSFM. Concretely, Dual-Spatial Mamba adopts a parallel strategy “Part-based + Whole-based”, ensuring that the model can capture fine-grained joint details, while preserving coarse-grained whole-body coordination. Meanwhile, the Bi-Temporal Mamba block employs a bidirectional selected SSMs to scan the time dimension both forward and backward, capturing both short-term cues and long-term dependencies to ensure temporal coordination. Finally, rather than simply stacking spatiotemporal modules, we propose DSFM, which removes redundant temporal features from Bi-Temporal Mamba, extracts complementary information, and fuses it with the spatial output from Dual-Spatial Mamba for more efficient spatiotemporal fusion. Thus, HiSTF Mamba accommodates both global and local, long-range and short-range, as well as spatial and temporal dimensions, thereby significantly improving the fidelity and realism of text-to-motion generation.
We conducted extensive experiments on the HumanML3D [20] and KIT [46] datasets, achieving excellent results across multiple key evaluation metrics. The experimental findings indicate that the proposed HiSTF Mamba can generate motion sequences with a high degree of semantic alignment to the text, naturally smooth transitions, and enhanced visual realism. The results on the HumanML3D dataset are particularly remarkable: HiSTF Mamba demonstrates significant improvements in various evaluation metrics, reducing the FID from 0.281 to 0.189 compared to the current state-of-the-art method, an almost 30% decrease, while maintaining leading performance in other key metrics.
Overall, our contributions to the text-to-motion community can be summarized as follows:
  1. 1.
    In the Dual-Spatial Mamba block, we propose a “Part-based + Whole-based” parallel strategy that captures both whole-body coordination and fine-grained joint-level details, thereby improving the model’s ability to represent complex human body structures.
  2. 2.
    We design a Dynamic Spatiotemporal Fusion Module (DSFM) to perform redundancy removal and extraction of complementary information for temporal features, which is then fused with spatial features to construct a more efficient spatio-temporal representation. This significantly enhances the realism and smoothness of the generated motion sequences.
  3. 3.
    Extensive experiments on HumanML3D demonstrate that HiSTF Mamba achieves outstanding results across multiple metrics. Notably, it reduces the FID score from 0.283 to 0.189, a reduction of nearly 30%, while also achieving leading performance on other metrics. This validates HiSTF Mamba’s superiority in balancing high fidelity and strong text relevance.
2 Related Work
2.1 Text-to-Motion Generation
Early text-to-motion works could only generate limited motions, such as Text2Action [1], which is restricted to synthesizing upper-body motion sequences from text descriptions by combining Recurrent Neural Network (RNN) and GAN. Meanwhile, JL2P [2] achieves a constrained one-to-one mapping between text and motion by learning a joint embedding space for language and pose. To address the problem that a text can only correspond to a determined motion, Guo et al. [20] proposed a motion generation model based on conditional VAE, which can stochastically generate human motions of variable lengths. TEMOS [43] utilizes a transformer-VAE to learn a joint latent space for both motion and text modalities, achieving alignment between motion and text within this latent space. TEACH [4] further extends this capability to longer motion sequences. Similarly, MotionCLIP [50] directly aligns the motion space with the CLIP space [48], effectively integrating the rich semantic knowledge of CLIP into the motion representation. Subsequently, VQ-VAE [52] is introduced in TM2T [21] to discretize motion sequences into tokens. This discrete representation facilitates the subsequent text-to-motion translation process. These VQ-VAE-based methods [57, 45, 22, 44] have since dominated the text-to-motion task, significantly outperforming other types of methods on the evaluation metric FID. This success is largely attributed to the discrete token representation, where each token inherently represents a brief snippet of real motion, significantly reducing the discrepancy between generated motions and real motions. However, the trade-off is that representing continuous motion data as a finite set of discrete tokens inevitably leads to information loss, thereby reducing the diversity of the generated motions.
In addition to GAN and VAE, the diffusion model [24], as a generative model, has also been widely adopted in the field of motion generation. Diffusion models have demonstrated their ability to produce high-quality and diverse samples while ensuring stable training, as evidenced in the realm of text-to-image [49]. MotionDiffuse [59] is the first framework to apply diffusion models to text-driven motion synthesis. MDM [51] employs a transformer encoder as its denoiser, directly predicting clean samples rather than noise. Due to the time-consuming and inefficient nature of conducting diffusion natively in the original motion data domain, MLD [9] opts to perform the diffusion process within the latent space of motion. Recently, MotionLCM [13] has introduced consistency distillation [40] on the foundation of MLD to tackle the inefficiency of diffusion sampling, further accelerating motion generation to real-time levels. Unlike previous methods, EMDM [62] proposes sampling motion from the diffusion model with fewer sampling steps during the denoising phase, achieving significant acceleration without compromising motion quality. Inspired by Mamba [16], Motion Mamba [61] introduces the Bidirectional Spatial Mamba block and the Bidirectional Temporal Mamba block into the U-net architecture, aiming to enhance the consistency of motion generation.
2.2 State Space Models
Transformers [53] excel in sequence modeling but often incur prohibitive quadratic complexity for high-resolution vision tasks. In response, State Space Models (SSMs)[18, 19] perform linear recurrences to capture long-range dependencies. For example, LSSL[19] employs continuous-time linear memory, while S4 [18], leveraging HiPPO [17] and FFT-based convolutions, significantly reduces computational overhead. Building on S4, Mamba [16] introduces dynamic gating and a parallel scan, accelerating inference without sacrificing global context.
In vision tasks, SSM-based techniques alleviate Transformers’ high computational cost by serializing images or videos into manageable sequences [16]. For example, Vision Mamba [63] and VMamba [37] replace self-attention with scanning strategies to achieve performance comparable to Transformers, whereas PlainMamba [56] and LocalMamba [25] adopt continuous or local scans to diminish spatial discontinuities and refine fine-grained segmentation. EfficientVMamba [42] incorporates dilated convolutions and stride sampling, maintaining accuracy under stringent resource constraints. Beyond scanning alone, additional modules can bolster local feature extraction: convolutional branches [7, 23, 65] are integrated into MambaIR [23] to merge convolutional and SSM layers for improved image restoration. Furthermore, MMA [10] and VL-Mamba [47] interleave attention-based components with SSMs to excel in super-resolution and multimodal fusion, while SpikeMba [35] and VideoMamba [34] leverage SSMs for video generation and temporal modeling, efficiently handling lengthy sequences and reducing computational costs.
3 Method
Figure 1: (Left) HiSTF Mamba overview. The text condition, the noise step $t$, and the noise-augmented sequence $\{x_{t}^{1},x_{t}^{2},\ldots,x_{t}^{N}\}$ are fed into the model. The text condition is encoded by the CLIP [48] text encoder, randomly masked, and projected into latent space to obtain the vector $c$. Next, $t$ is processed by an MLP and added to $c$ to form the token $Z_{\text{tk}}$. Finally, these inputs are used by the model to predict the denoised sequence $\{\hat{x}_{0}^{1},\hat{x}_{0}^{2},\hat{x}_{0}^{3},\ldots,\hat{x}_{0}^{N}\}$. (Right) Sample HiSTF Mamba. The text condition $c$, the noise step $t$, and a random noise $X_{T}$ are fed into the model to predict a clean sequence $\hat{x}_{0}$. The diffusion process then generates $x_{T-1}$ from $\hat{x}_{0}$. Repeating this procedure $T$ times yields a clean sequence $x_{0}$ that conforms to the text condition. DMP-Solver++ [39] is used to accelerate sampling.
In this section, we delineate the overall architecture and operational principles of the HiSTF Mamba, which aims to generate human motion guided by textual inputs. We begin by describing the theoretical foundations of Mamba [16], followed by an overview of the HiSTF Mamba. Subsequently, we detail the spatiotemporal modules: Dual-Spatial Mamba and Bi-Temporal Mamba, then proceed to discuss DSFM, which effectively fuses these spatiotemporal features. Finally, we introduce the loss function design for HiSTF Mamba.
3.1 Preliminaries
In S4  [18], under a discrete-time setting, a linear time-invariant (LTI) state space model is:
$\begin{cases}h_{t}=A\,h_{t-1}+B\,x_{t},\\[6.0pt] y_{t}=C\,h_{t},\end{cases}$ (1)
where $h_{t}\in\mathbb{R}^{N}$ is the hidden state, and $x_{t},y_{t}\in\mathbb{R}$ are the input and output (extendable to multi-channel). The parameters $A,B,C$ are trainable but, as an LTI system, cannot selectively emphasize or suppress specific inputs.
Selective SSMs [16] address this by making some key parameters vary dynamically with the input sequence, thus providing “content selectivity”. Specifically, some parameters, such as $B_{t}$, $C_{t}$, and $\Delta_{t}$, are determined by the current input $x_{t}$. Concretely:
$\begin{cases}h_{t}=A(\Delta_{t})\,h_{t-1}+B_{t}\,x_{t},\\[6.0pt] y_{t}=C_{t}\,h_{t},\end{cases}\quad\begin{cases}B_{t}&=s_{B}\bigl{(}x_{t}\bigr{)},\\ C_{t}&=s_{C}\bigl{(}x_{t}\bigr{)},\\ \Delta_{t}&=\tau_{\Delta}\!\bigl{(}W_{\Delta}\,x_{t}\bigr{)},\end{cases}$ (2)
By discretizing $A,B$ into time-varying matrices $A(\Delta_{t}),B_{t}$, the system becomes time-varying, allowing it to filter out irrelevant information. This breaks LTI assumptions, preventing direct use of FFT-based global convolution. However, if $A(\Delta_{t})$ is diagonal or similarly structured, a parallel scan on GPUs can achieve near $\mathcal{O}(L)$ complexity, and fusion or recomputation can reduce memory overhead.
3.2 HiSTF Mamba
Figure 2: (Left) Bi-Temporal Mamba. The input sequence is normalized, and then processed forward and backward to capture richer temporal dependencies for smooth, coherent motion. (Right) Dual-Spatial Mamba. The Part-based branch splits the full-body motion into six parts, extracts fine-grained local features via a 2D dynamic convolution and Part SSMs (Part Mamba), and recombines them in Part2Whole. Meanwhile, the Whole-based branch models the entire body (Whole SSMs) to learn globally coordinated patterns. Finally, local and global features are summed, obtaining a representation that is both well-coordinated and captures joint-level details.
Existing methods face two main issues. First, when the entire motion matrix is fed into the network at once, earlier joint features become overshadowed, leading to the loss of fine-grained details. Second, stacking spatio-temporal modules not only fails to effectively learn and fuse spatiotemporal features, resulting in duplicate or overlapping feature extraction that increases computational overhead and reduced fidelity. To address these problems, we propose HiSTF Mamba, as illustrated in Fig. 1 a hierarchical spatiotemporal fusion Mamba method for text-to-motion generation.
HiSTF Mamba addresses the difficulty of extracting fine-granularity joint details and the inability to effectively fuse spatio-temporal features through three specialized components: N Bi-Temporal Mamba blocks, one Dual-Spatial Mamba block, and a DSFM module. By modeling part-based and whole-based motions in parallel with Dual-Spatial Mamba and then dynamically fusing them with temporal features learned by Bi-Temporal Mamba using the DSFM module, HiSTF Mamba ensures whole-body coordination while preserving fine-grained joint details and efficiently fusing temporal features.
Our proposed HiSTF Mamba is built upon MDM [51]’s diffusion and sampling processes, and we integrate a DMP-Solver++ [39] strategy to accelerate sampling, reducing the original 1000 steps to 10 or 15. $c$, Specifically, HiSTF Mamba takes a processed text condition a noise step $t$, and a motion sequence $\{x_{t}^{1},x_{t}^{2},\dots,x_{t}^{N}\}$ subjected to $t$ diffusion steps, and outputs the corresponding clean sequence $\{x_{0}^{1},x_{0}^{2},\dots,x_{0}^{N}\}$. This design balances global motion patterns with fine-grained joint dynamics, producing more accurate reconstructions. Here, $x_{0}^{i}\in\mathbb{R}^{T\times D}$, where $T$ denotes the sequence length and $D$ is the dimensionality of each joint representation.
3.3 Bi-Temporal Mamba and Dual-Spatial Mamba
As shown in Figure 2, Bi-Temporal Mamba employs a bidirectional scanning mechanism to capture the contextual information of motion sequences in both forward and backward directions, thereby generating more coordinated and coherent motion sequences. Dual-Spatial Mamba adopts parallel Part-based and Whole-based learning, effectively capturing motion features that exhibit both fine-grained joint details and whole-body coordination.
Bi-Temporal Mamba. To generate coherent and smooth motion sequences, we employ Selected SSMs featuring a bidirectional scanning mechanism. We begin with an input sequence $T_{l-1}\in\mathbb{R}^{(B\times T\times D)}$, where $B$ is the batch size, $T$ is the sequence length, and $D$ is the feature dimension. First, $T_{l-1}$ is normalized and subsequently projected by a linear layer into a new representation space $\mathbb{R}^{(B\times T\times E)}$, yielding two transformed sequences $X$ and $Z$. We then perform forward and backward scans over $X$, each governed by learnable parameters $\Delta$, $B$, and $C$, which dynamically adjust the state update equation following Equation (2)
The forward and backward outputs are then elementwise multiplied with the nonlinear activation $\mathrm{SiLU}(Z)$ and fused. Finally, a linear projection maps the fused representation back to the original dimension $\mathbb{R}^{(B\times T\times D)}$.This bidirectional scanning strategy is particularly effective for modeling motion sequences. The forward branch captures how earlier actions influence and shape the current state, while the backward branch leverages future actions as contextual cues. By combining these two directions within a temporal recursion framework, the model more thoroughly encodes transitions and dependencies across time steps, ultimately producing more coordinated and smooth motion sequences.
Figure 3: DSFM: Redundant information is removed and complementary information is extracted from the temporal features produced by the N Temporal Mamba blocks, yielding the refined temporal feature $\mathbf{T}$. This feature is then fused with the spatial feature $\mathbf{S}$ from the Space Mamba block to obtain a more expressive spatiotemporal feature representation – fused feature.
Dual-Spatial Mamba. To address the limitation that existing spatial modules only focus on global motion features while neglecting fine-grained joint-level details, we propose the Dual-Spatial Mamba. This module performs two types of learning at the same time through Multi-Granular Body-Spatial Modelling, as described in Figure 2: on the Part-based the entire motion matrix is decomposed into 6 body part matrices, so that each input is limited to body part joints, allowing the model to focus on learning body part joint details, avoiding too much joint information being input at the same time, which can be severely compressed early on. The Whole-based retains the entire motion matrix inputs to learn whole-body motion coordination. Thus, Dual-Spatial Mamba adopts Part-based and Whole-based parallel learning to capture both local and global motion features, which are then fused via an element-wise addition strategy.
Specifically, the input sequence is normalized. In the Part-based learning phase, we first map the normalized full-body motion sequence back to its original dimensionality. Drawing upon insights from ParCo [64], we design a Whole2Part module to split the whole-body motion into six parts. R.Leg, L.Leg, Backbone, Root, R.Arm, and L.Arm. This design follows the common practice of 3D human body models (e.g., SMPL, MMM), where Kinematic Trees typically represent the skeleton as five main chains (the four limbs plus the backbone); ParCo adds a separate Root component to account for global trajectories, thereby leading to six sub-motions in total. Formally, given a motion sequence $x=\{x^{1},x^{2},x^{3},\dots,x^{N}\}$, it is partitioned into $p_{i}=\{p_{i}^{1},p_{i}^{2},p_{i}^{3},\dots,p_{i}^{N}\},\quad i\in\{1,2,\ldots,S\}$, where $S$ denotes the total number of body parts. After segmenting the motion into six parts, the features of each part are projected to a latent dimension and fed into a 2D dynamic convolution network. To better capture the distinct motion patterns of each part, six learnable convolution kernels are used, each corresponding to a specific body part, thus improving the model’s capability to capture fine-grained local details. Next, we stack the outputs of these six parts along the batch dimension and feed them into the Selected SSMs (Part Mamba) module, which further models local dependencies and outputs refined local motion features. Finally, through the Part2Whole process, we restore the local characteristics to a holistic representation of the entire body, thus obtaining fine-grained characteristics at the joint level. In Whole-based learning phase, the full-body motion matrix is fed into a 1D convolutional network to extract key motion features of different parts. Subsequently, by leveraging the Selected SSMs (Whole SSMs) module’s advantages in sequence processing and long-range dependency modeling, inter-joint relationships and latent motion patterns at the global level are captured, thereby yielding global motion features. Finally, Part-based and Whole-based motion features are summed element-wise to obtain a representation that combines both global motion patterns and fine-grained joint-level detail features.
3.4 Dynamic Spatiotemporal Fusion Module
To better integrate temporal and spatial features and enable the model to learn more expressive spatiotemporal representations, Inspired by the DMCA module in Sparx [38], we propose a dynamic spatiotemporal fusion module (DSFM). Differently, DSFM not only reduces redundancy and extracts complementary features among the output features from the $N$ layers of Temporal Mamba, but also modifies the fusion strategy to effectively integrate this refined temporal features with spacial features to obtain a expressive spatiotemporal representation. Specifically, as shown in Fig. 3, let $Y\in\mathbb{R}^{(C\times T\times D)}$ represent the output features of $N$ Temporal Mamba layers, where $Y=\{Y_{1},Y_{2},\dots,Y_{N}\}$. In practice, the output of Temporal Mamba is originally in the form $Y\in\mathbb{R}^{(T\times D)}$. To facilitate an effective division into $Y_{q}$ and $Y_{k}$ in the subsequent processing steps, we apply an unsqueeze operation to reshape it into $Y\in\mathbb{R}^{(C\times T\times D)}$. Additionally, let $S\in\mathbb{R}^{(T\times D)}$ denote the output features of Space Mamba. Here, $C$ represents the number of channels, while $T$ and $D$ denote the sequence length and the feature dimension respectively.
To obtain refined temporal information, the first $N-1$ timporal features $\{Y_{1},Y_{2},\dots,Y_{N-1}\}$ are stacked along the channel dimension and then projected into $\mathbb{R}^{(2C\times T\times D)}$. The projected features are subsequently split into two components, each squeezed to $\mathbb{R}^{(T\times D)}$. One component serves as the query $Y_{q}$, while the other serves as the key $Y_{k}$.
The attention matrix is computed using $Y_{v}$ and $Y_{k}$. According to [38], this attention matrix is independent of the spatial dimension of $Y_{v}$ and $Y_{k}$. Hence, a reduction by a factor of $s$ can be applied, mapping $Y_{k},Y_{v}$ to $\mathbb{R}^{(T\times\tfrac{D}{s})}$. This step considerably lowers computational overhead and memory usage while preserving key temporal relationships. The resulting matrix is then used to weight the query $Y_{q}$, yielding the refined temporal feature.
Next, the refined temporal features are integrated with the spatial features. The resulting representations are concatenated along the feature dimension, forming $\mathbb{R}^{(T\times 2D)}$. Finally, a linear projection is applied to map the concatenated features back to $\mathbb{R}^{(T\times D)}$, ensuring a unified representation. This fusion mechanism allows the model to effectively capture complementary spatiotemporal patterns, enhancing its ability to learn robust feature representations.
Mathematically, the Dynamic Spatiotemporal Fusion Module (DSFM) can be formulated as follows:
$Y_{k},Y_{q}$ $=\text{squeeze}\!\Bigl{(}\text{split}\bigl{(}W_{1}\cdot\text{concat}(Y_{1},Y_{2},\dots,Y_{N-1})\bigr{)}\Bigr{)},$ (3)
$V$ $=R\bigl{(}(Y_{N})^{T}\bigr{)},\quad K=R\bigl{(}(Y_{k})^{T}\bigr{)},\quad Q=Y_{q},$
$\mathbf{T}$ $=\frac{\text{Softmax}(QK^{T})}{\sqrt{N/r}}\,V,$
$Z$ $=W_{2}\cdot\text{Concat}\bigl{(}S,\;T\bigr{)},$
where $W$ represents a linear projection. Specifically, $W_{1}\in\mathbb{R}^{((N-1)C\times 2C)}$ performs a projection along the channel dimension, while $W_{2}\in\mathbb{R}^{(2D\times D)}$ projects along the feature dimension. The function $R$ denotes a reduction operation that decreases the feature dimension, implemented via a strided convolution. Additionally, $K,V\in\mathbb{R}^{(T\times\tfrac{D}{s})}$ represent the key and value matrices after dimensionality reduction.
Table 1: Comparison of results on the HumanML3D test set. The right arrow (→) indicates that values closer to real motions correspond to better performance. Each evaluation was run 20 times, and “±” denotes the 95% confidence interval. Bold text highlights the best scores, while underlined entries mark the second-best.
Methods R Precision↑ FID↓ MM Dist↓ Diversity→ MModality↑
Top 1 Top 2 Top 3
Real 0.511±.003 0.703±.003 0.797±.002 0.002±.000 2.974±.008 9.503±.065 -
T2M  [53] 0.457±.002 0.639±.003 0.740±.003 1.067±.002 3.340±.008 9.188±.002 2.090±.083
TEMOS  [43] 0.424±.002 0.612±.002 0.722±.002 3.734±.028 3.703±.008 8.973±.071 0.368±.018
MDM  [51] 0.320±.005 0.498±.004 0.611±.007 0.544±.044 5.566±.027 9.559±.086 2.799±.072
MotionDiffuse  [59] 0.491±.001 0.681±.001 0.782±.001 0.630±.001 3.113±.001 9.410±.049 1.553±.042
MLD  [9] 0.481±.003 0.673±.003 0.772±.002 0.473±.013 3.196±.010 9.724±.082 2.413±.079
MotionGPT  [60] 0.364±.005 0.533±.003 0.629±.004 0.805±.002 3.914±.013 9.972±.026 2.473±.041
M2DM  [30] 0.497±.003 0.682±.002 0.763±.003 0.352±.005 3.134±.010 9.926±.073 3.587±.072
MMDM  [8] 0.435±.006 0.627±.006 0.733±.007 0.285±.032 3.363±.029 9.398±.088 2.701±.083
Motion Mamba  [61] 0.502±.003 0.693±.002 0.792±.002 0.281±.009 3.060±.058 9.871±.084 2.294±.058
HiSTF Mamba (10-steps) 0.488±.005 0.685±.004 0.784±.005 0.189±.018 3.101±.022 9.712±.090 2.529±.044
HiSTF Mamba (15-steps) 0.504±.005 0.699±.005 0.798±.005 0.249±.023 3.053±.022 9.383±.091 2.276±.036
Table 2: Comparison of quantitative results on the KIT-ML dataset
Methods R Precision↑ FID↓ MM Dist↓ Diversity→ MModality↑
Top 1 Top 2 Top 3
Real 0.424±.005 0.649±.006 0.779±.006 0.031±.004 2.788±.012 11.08±.097 -
T2M  [53] 0.370±.005 0.569±.007 0.693±.007 2.770±.109 3.401±.008 10.91±.119 1.482±.065
MDM  [51] 0.164±.004 0.291±.004 0.396±.004 0.497±.021 9.191±.022 10.85±.109 1.907±.214
MotionDiffuse  [59] 0.417±.004 0.621±.004 0.739±.004 1.954±.062 2.958±.005 11.10±.143 0.730±.013
MLD  [9] 0.390±.008 0.609±.008 0.734±.007 0.404±.027 3.204±.027 10.80±.117 2.192±.071
M2DM  [30] 0.416±.004 0.628±.004 0.743±.004 0.515±.029 3.015±.017 11.42±.097 3.325±.037
MMDM  [8] 0.386±.007 0.603±.006 0.729±.006 0.408±.022 3.215±.026 10.53±.100 2.261±.144
Motion Mamba  [61] 0.419±.006 0.645±.005 0.765±.006 0.307±.041 3.021±.025 11.02±.098 1.678±.064
HiSTF Mamba (10-steps) 0.437±.006 0.651±.006 0.772±.006 0.289±.021 2.846±.018 10.92±.096 1.512±.088
HiSTF Mamba (15-steps) 0.440±.006 0.657±.006 0.774±.006 0.293±.017 2.819±.015 10.93±.099 1.347±.056
3.5 Loss
The primary training loss function of HiSTF Mamba is defined as $\mathcal{L}_{\text{simple}}$, where in HiSTF Mamba we predict the simple objective directly, i.e., $\hat{x}_{0}=f(x_{t},t,c)$. the primary loss is defined as:
$\mathcal{L}_{\text{simple}}=\mathbb{E}_{x_{0}\sim q(x_{0}\mid c),\,t\sim[1,T]}\Bigl{[}\bigl{\|}x_{0}-f(x_{t},t,c)\bigr{\|}\Bigr{]}.$ (4)
In addition, we incorporate the position loss, foot contact loss, and velocity loss to encourage physically plausible motions:
$\mathcal{L}_{\text{pos}}=\frac{1}{N}\sum_{i=1}^{N}\left\lVert FK\bigl{(}x_{0}^{i}\bigr{)}-FK\bigl{(}\hat{x}_{0}^{i}\bigr{)}\right\rVert_{2}^{2},$ (5)
$\mathcal{L}_{\text{foot}}=\frac{1}{N-1}\sum_{i=1}^{N-1}\left\lVert\bigl{(}FK(x_{0}^{i+1})-FK(x_{0}^{i})\bigr{)}\cdot f_{i}\right\rVert_{2}^{2},$ (6)
$\mathcal{L}_{\text{vel}}=\frac{1}{N-1}\sum_{i=1}^{N-1}\left\lVert\bigl{(}(x_{0}^{i+1}-x_{0}^{i})-(\hat{x}_{0}^{i+1}-\hat{x}_{0}^{i})\bigr{)}\right\rVert_{2}^{2}.$ (7)
Thus, the total loss is formulated as:
$\mathcal{L}=\mathcal{L}_{\text{simple}}+\lambda_{\text{pos}}\mathcal{L}_{\text{pos}}+\lambda_{\text{vel}}\mathcal{L}_{\text{vel}}+\lambda_{\text{foot}}\mathcal{L}_{\text{foot}}.$ (8)
4 Experiments
4.1 Datasets and Evaluation Metrics
To evaluate the effectiveness of our method, we conduct experiments on two widely-used text-to-motion datasets: HumanML3D [20] and KIT-ML [46]. HumanML3D consists of 14,616 motion sequences and 44,970 textual descriptions. Meanwhile, KIT-ML, introduced by Plappert et al., comprises 3,911 motion sequences and 6,278 natural language annotations. To ensure a fair comparison, we follow the evaluation protocol established in [20] for subsequent experiments.
Five evaluation metrics from [20] are adopted, including (1) Frechet Inception Distance (FID), which serves as the primary metric to assess the difference in feature distributions between generated motions and real motions; (2) R-Precision evaluates how well a generated motion aligns with its corresponding textual description. It measures the retrieval accuracy by checking whether the ground truth description ranks among the top-k closest matches based on Euclidean distances between motion and text features; (3) MultiModal Distance (MM Dist), which calculates the average distance between generated motions and their corresponding text descriptions; (4) Diversity measures the variety of generated motions and is computed as the variance of motion features; (5) Multimodality evaluates the variety of motions generated from the same text description.
4.2 Comparative Studies
We evaluated our method on the HumanML3D [20] and KIT-ML [46] datasets and compared it with several existing state-of-the-art approaches. Specifically, we adopted the network structure described in Section 3.2, setting the number of model layers to $N=3$ and $L=2$. Meanwhile, to accelerate the sampling process, we employed DMP-solver++ and selected $step=10$ and $step=15$ for inference. As shown in Tables 1 and 2, our method achieves significant performance improvements on both the HumanML3D and KIT-ML datasets. On HumanML3D, when $step=10$ is used, our FID metric is further reduced from 0.281 (obtained by the state-of-the-art Motion Mamba) to 0.189, representing an improvement of more than 30%, and we also observe increased R-Precision and MultiModal Distance. On KIT-ML, our method similarly outperforms existing approaches in terms of R-Precision, FID, and MM Dist, with the MM Distance reduced to 2.819.
4.3 Ablation Studies
We investigate how the number of Bi-Temporal Mamba blocks ($N$), the total number of layers ($L$), and the presence of key components: DSFM and part-based learning in Dual-Spatial Mamba) affect model performance. Table 3 provides the results on the HumanML3D dataset with an inference step of 10. The complete HiSTF Mamba configuration, which includes parallel Part-based and Whole-based learning in Dual-Spatial Mamba, Bi-Temporal Mamba, the DSFM module, and uses $N=3$ with $L=2$, achieves the best FID of 0.189. By contrast, merely increasing $N$ to 4 or $L$ to 3 does not consistently improve performance; the FID ranges from 0.303 to 0.397. This finding indicates that simply increasing the number of Bi-Temporal Mamba modules or stacking more spatio-temporal layers, the FID metric does not steadily decrease; instead, it increases or fluctuates within a certain range. This observation indicates that merely stacking spatiotemporal modules does not effectively improve generation quality. In contrast, removing the DSFM module leads to a significant increase in FID form 0.189 to 0.319, showing that the effective fusion of spatiotemporal features provided by DSFM is crucial for achieving higher-quality motion generation. e part-based parallel strategy from Dual-Spatial Mamba and omitting the DSFM module. Discarding part-based learning raises the FID to 0.349, while removing DSFM increases it to 0.319.Furthermore, eliminating part-based modeling in Dual-Spatial Mamba also increases the FID from 0.189 to 0.349, suggesting that losing the ability to learn fine-grained joint details reduces the quality of the generated motions.
In summary, the performance improvement of HiSTF Mamba primarily arises from the Dual-Spatial Mamba’s local-global parallel spatial modeling and the DSFM module’s dynamic fusion of spatiotemporal features. These two components work in synergy to maintain overall motion coherence, capture the fine-grained details of human joints, effectively fuse spatiotemporal features, and ultimately achieve a significantly lower FID.
Table 3: Ablation study on HumanML3D with inference step=10. We vary $N$, $L$, remove part-based learning in Dual-Spatial Mamba, and remove the DSFM module. For simplicity, we use BTM to represent Bi-Temporal Mamba, and P-b and W-b to denote the part-based learning and whole-based learning in Dual-Spatial Mamba, respectively.
BTM W-b P-b DSFM N L FID
2 2 0.313
3 2 0.189
4 2 0.303
2 3 0.335
3 3 0.397
4 3 0.363
3 2 0.349
3 2 0.319
4.4 Implementation Details
All models in the HiSTF Mamba use the AdamW optimizer, with the learning rate maintained at $1\times 10^{-4}$. The average training step is about 570k, and we set the batch size to 64. We fix the diffusion step count at 1000 and use 10 or 15 sampling steps. On a single NVIDIA GEFORCE RTX 4090 GPU, training on the HumanML3D dataset takes approximately 37.5 hours. The average inference time is about 0.69 seconds.
5 Conclusion
In this paper, we have introduced HiSTF Mamba, a novel method for text-to-motion generation that integrates local and global spatial features with bidirectional temporal modeling. Our Dual-Spatial Mamba block effectively captures fine-grained joint dynamics alongside coarse whole-body movements, while the Bi-Temporal Mamba block encodes both short-term and long-term dependencies in a bidirectional manner. We also introduced the DSFM block to perform dynamic fusion of refined temporal features with spatial representations, thereby reducing redundancy and enhancing overall spatiotemporal expressiveness.
Extensive experiments on HumanML3D and KIT-ML demonstrate that HiSTF Mamba not only reduces FID by nearly 30% in HumanML3D relative to state-of-the-art methods, but also delivers high-fidelity and semantically aligned motion sequences. In addition, our method achieves these gains with efficient training times and rapid inference. We believe HiSTF Mamba offers a promising approach for generating more realistic, detail-rich human motions, and serves as a foundation for future work in text-to-motion generation.